The motor on a construction-grade hydraulic shovel exerts 3.11 × 107 Pa of pressure on a fluid tank. The fluid tank is connected to a piston that has an area of 153 cm2. How much force does the piston exert?
1
Expert's answer
2021-02-08T18:39:06-0500

"\text{according to Pascal's law}"

"p=\frac{F_1}{S_1}=\frac{F_2}{S_2}"

"F_1=p*S_1"

"p = 3.11*10^7Pa"

"S_1=153cm^2=153*10^{-4}m^2"

"F_1=3.11*10^7*153*10^{-4}=475830N"

Answer:475830 N